1. 21) Find the value of 20.10 Provide a sketch of the solution. area =0.10 Zo.10 Z
how do I solve this problem? please explain how to solve it using the calculator. thanks!

2. 23) 3) A binomial A binomial probability with parameters, n = 9 and ps a) Find the probabilities and construct a binomial pro
how do I solve part b? thanks!
Answer #1

Ans:

1)

z0.1 indicates the z score which has area 0.1 to left of  the z score.

P(Z<=z)=0.1

z=normsinv(0.1)

z=-1.282

(use InvNorm function of calculator)

224 2=-3 2-2 2-1 230 231 232 233 234

2)

b)

Binomial distribution with n=9 and p=0.8

mean=np=9*0.8=7.2

standard deviation=sqrt(9*0.8*(1-0.8))=1.2